### v2

# EvaGaussians: Event Stream Assisted Gaussian Splatting from Blurry Images

Abstract
3D Gaussian Splatting (3D-GS) has demonstrated exceptional capabilities in synthesizing novel views of 3D scenes. However, its training is heavily reliant on high-quality images and precise camera poses. Meeting these criteria can be challenging in non-ideal real-world conditions, where motion-blurred images frequently occur due to high-speed camera movements or low-light environments. To address these challenges, we introduce Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel approach that harnesses event streams captured by event cameras to facilitate the learning of high-quality 3D-GS from blurred images. Capitalizing on the high temporal resolution and dynamic range offered by event streams, we seamlessly integrate them into the initialization and optimization of 3D-GS, thereby enhancing the acquisition of high-fidelity novel views with intricate texture details. To remedy the absence of evaluation benchmarks incorporating both event streams and RGB frames, we present two novel datasets comprising RGB frames, event streams, and corresponding camera parameters, featuring a wide variety of scenes and various camera motions. We then conduct a thorough evaluation of our method, comparing it with leading techniques on the provided benchmark. The comparison results reveal that our approach not only excels in generating high-fidelity novel views, but also offers faster training and inference speeds. Video results are available at the project page.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author.
1 Introduction
Figure 1: EvaGaussians integrates blurry images and event streams to reconstruct sharp 3D-GS for novel view synthesis.
Novel view synthesis from 2D image collections has presented a persistent challenge within the field of computer vision and computer graphics. This task stands as a fundamental component in various vision applications, such as virtual reality [45, 16, 48, 36], robotics navigation [32, 54, 46], scene understanding [14, 19, 20], and many others, thereby prompting significant research efforts over the last decades. Amid pioneering works,
3D Gaussian Splatting (3D-GS) [13] achieves notable success in generating high-fidelity novel views. It learns 3D Gaussians with the lightweight learnable parameters, and leverages a tile-based rasterization technique to render novel views, thereby surpassing NeRFs [26] in both training and rendering efficiency. However, the optimization of 3D-GS heavily relies on accurate camera poses and point cloud initialization produced by COLMAP [34], which necessitates high-quality images without blurring and with adequate lighting. Fulfilling such conditions can be challenging in real-world situations. For example, in UAVs and robotics, rapid camera movement is common when capturing images or recording videos, which often result in significant motion blur. The mismatched features between blurred images can lead to inaccurate pose calibrations and point cloud initialization, thereby hindering the training process of 3D-GS.
Recent studies have demonstrated the significant potential of event-based cameras in alleviating motion blur in images captured by conventional frame-based cameras [29, 11, 18, 37, 49, 12, 36]. Serving as an innovative bio-inspired visual sensor, event cameras asynchronously report the logarithmic intensity changes of each pixel captured, and can record higher temporal resolution and dynamic range data in contrast to conventional cameras. Motivated by this, prior works [30, 2], have attempted to leverage the event streams captured by event cameras to supervise the training of NeRFs. However, achieving real-time rendering and synthesizing high-fidelity novel views with intricate details poses substantial challenges for these methods.
To address these challenges, we introduce Event Stream Assisted Gaussian Splatting (EvaGaussians), which leverages the event streams captured by event cameras to enhance the learning of high-quality 3D-GS from motion-blurred images. Harnessing the exceptional temporal resolution and dynamic range offered by event streams, we use them to assist in the initialization of 3D-GS, and incorporate them to jointly optimize 3D-GS and camera trajectories of blurry images through a blur reconstruction loss and an event reconstruction loss. Due to the geometric ambiguity caused by blurry images, we further propose two event-assisted depth regularization terms to stabilize the geometry of 3D-GS. Through optimizing the 3D-GS in a progressive manner, our method can recover a high-quality 3D-GS that facilitates the real-time generation of high-fidelity novel views. To summarize, our contributions can be delineated as follows:
  • We propose Event Stream Assisted Gaussian Splatting (EvaGaussians), a framework tailored for reconstructing a high-quality 3D-GS from motion-blurred images with the assistance of event camera. Once trained, our method is capable of recovering intricate details of the input blurry images and allows high-fidelity real-time novel view synthesis.
  • We contribute two novel datasets, including a synthetic dataset containing diverse scenes with various scales, and a real-world dataset captured by the color DAVIS346 event camera [1], both feature various camera motions. We believe they will set a benchmark for future researches.
  • We conduct a comprehensive evaluation of the proposed method and compare it with several strong baselines. The results reveal that our approach not only excels in generating high-fidelity novel views but also provides faster training and inference speeds.
2 Related Works
2.1 Reconstructing 3D Scene from Blurry Images
Reconstructing a high-quality 3D Scene typically requires high-fidelity, sharp images as supervision. However, motion-blurred images often occur in real world scenarios, thus hindering accurate reconstruction of 3D scenes. Several studies have been proposed to address this issue. For example, Deblur-NeRF [24] and DP-NeRF [17] attempted to learn a blur formation kernel to model the image blurring process. BAD-NeRF [41] further physically modeled the blurry images formation process, and adopted a bundle-adjustment strategy to jointly optimize NeRF parameters and the camera poses during the exposure time. These NeRF-based methods lacked real-time rendering capabilities and suffered from extended training times. With the rapid advancement of 3D-GS, a concurrent work, BAD-Gaussians [52], proposed to utilize 3D-GS as representation and follow the blur modeling and bundle-adjustment strategy adopted in [41] to achieve deblurring reconstruction. Although it achieved real-time rendering and faster convergence compared with prior works, it still struggled to handle severely blurred images in which COLMAP [34] will fail to produce the initial point clouds. Furthermore, it employed linear interpolation between the start and end camera poses to model camera trajectory during exposure time, necessitating careful selection of poses for more stable optimization.
2.2 Reconstructing 3D Scene from Event Streams
Motivated by the exceptional properties offered by event cameras, several studies attempted to reconstruct 3D scenes from event streams captured by event cameras, particularly in low-light conditions with fast camera motion. For example, EventNeRF [33], Ev-NeRF [10] and other concurrent works [43, 44, 47, 51, 48] explored the reconstruction of a 3D representation from a rapidly moving event camera. Robust e-NeRF [22] and its variants [23] further extended this task to the more challenging scenario of non-uniform camera motion, taking into account the refractory period of event cameras. These methods were typically designed to be supervised solely by information captured from a single event camera. Recently, E-NeRF [15], ${\textnormal{E}^{2}}$eRF [30], and EvDeblurNeRF [3] proposed to jointly utilize event streams captured by event cameras and motion-blurred images captured by standard frame-based cameras to reconstruct a NeRF representation. Compared to methods that rely solely on event cameras, these methods can recover accurate color details. Additionally, in contrast to RGB-only methods, they are better at handling motion blur. However, these NeRF-based methods suffer from long training and inference times, and face instability during training, which limit their further application.
Figure 2: Overview of EvaGaussians. We use event streams to assist in the initialization of 3D-GS and incorporate them to jointly optimize both 3D-GS and the camera trajectories of blurry images during the exposure time, utilizing a blur reconstruction loss and an event reconstruction loss. Additionally, we propose two event-assisted depth regularization terms to stabilize the geometry of 3D-GS.
3 Method
3.1 Preliminary
[…]
We can then represent the logarithmic intensity change as: $\mathbf{L}(t)-\mathbf{L}(s)=c\,\mathbf{E}(t)$, rewrite as $\mathbf{L}(t)=\mathbf{L}(s)+c\,\mathbf{E}(t)$, and subsequently obtain the actual intensity change:
$\mathbf{I}(t)=\mathbf{I}(s)\cdot\exp(c\,\mathbf{E}(t)).$ (2)
Therefore, when an image $\mathbf{I}(s)$ is captured at time $s$, and the event stream is recorded during the time interval $[s,t]$, the image $\mathbf{I}(t)$ can be obtained by warping $\mathbf{I}(s)$ using Eq. 2.
3.2 Event-assisted Initialization
The optimization of 3D-GS requires camera calibration and point cloud initialization using COLMAP [34]. However, this process can fail when dealing with images that have significant motion blur. Motion-blurred images are resulted from camera movements during the exposure time, which can be mathematically represented as:
[…]
where $\mathbf{B}$ denotes a captured blurry image, which is equivalent to averaging the instantaneous latent images $\mathbf{I}(t)$ during the exposure time $[s-\tau/2,s+\tau/2]$.
To obtain initial camera poses and point clouds for 3D-GS optimization, we first preprocess the motion-blurred images using the Event-based Double Integral (EDI) [29] model, which can be derived through substituting Eq. 2 into Eq. 3:
[…]
Given the predefined threshold $c$, a blurry image $\mathbf{B}$, and the recorded event stream $\mathbf{E}(t)$, the EDI model (Eq. 4) allows the derivation of $\mathbf{I}(s)$, following which the latent image $\mathbf{I}(t)$ at any moment within the exposure time can be estimated through Eq. 2. As shown in Figure. 2(A), given a total of $K$ blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$, for each of them, we uniformly sample $n$ time stamps during their exposure time to obtain a series of EDI-estimated latent images rich in texture features, denoted as $\{\mathbf{I}_{i}\}_{i=1}^{n}$, then obtain their poses $\{\mathbf{P}_{i}\}_{i=1}^{n}$ and the initial point cloud of the scene using COLMAP [34].
After initialization, a straightforward approach to optimizing the 3D-GS is to use the EDI-estimated latent images and poses as supervision. However, although these images provide more texture features than the original blurry image, they still do not fully recover the ideal latent image and exhibit relatively low visual quality, which also introduces inaccuracies into the camera poses, thereby leading to unsatisfactory optimization results. To more robustly recover a sharp 3D-GS from motion-blurred images, we propose to harness the advantages of event streams and seamlessly integrate them into the optimization process of 3D-GS.
3.3 Event-assisted Bundle Adjustment
As introduced in Eq. 3, during the exposure time, a motion-blurred image can be decomposed into a series of latent images along a specific camera trajectory, which can be roughly approximated by the EDI-produced camera poses $\{\mathbf{P}_{i}\}_{i=1}^{n}$ according to Eq. 4. Motivated by this, we jointly optimize these camera poses and the 3D-GS attributes in a bundle adjustment manner [41] to simultaneously recover the blur-formation camera trajectories and a sharp 3D-GS. As shown in Figure. 2(B), we add each of the EDI-produced camera poses a learnable offset $\{\mathbf{d}_{i}\}_{i=1}^{n}$ as correction parameters, resulting a learnable camera trajectory $\{\mathbf{\widetilde{P}}_{i}\}_{i=1}^{n}$, where $\mathbf{\widetilde{P}}_{i}=\mathbf{P}_{i}+\mathbf{d}_{i}$. In each training iteration, we simultaneously render $n$ images $\{\mathbf{\widetilde{I}}_{i}\}_{i=1}^{n}$ from the 3D-GS along the camera trajectory of the blurry view, and simulate the formation of motion-blurred images using a discrete approximation of Eq 3, expressed as $\mathbf{\widetilde{B}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{\widetilde{I}}_{i}$. Consequently, for a total of $K$ real-captured blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$, we can obtain their simulated versions $\{\mathbf{\widetilde{B}}^{j}\}_{j=1}^{K}$ through each corresponding learnable camera trajectory.
Blur Reconstruction Loss. With the simulated blurry images, we use the real captured blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$ to serve as image level supervision. Specifically, for each blurry image $\mathbf{B}^{j}$ and its simulated version $\mathbf{\widetilde{B}}^{j}$, we employ a blur reconstruction loss to minimize their photometric error, expressed as
$\mathcal{L}_{blur}=(1-\lambda_{1})\cdot\|\mathbf{B}^{j}-\mathbf{\widetilde{B}}^{j}\|_{1}+\lambda_{1}\cdot\text{D-SSIM}(\mathbf{B}^{j},\mathbf{\widetilde{B}}^{j}).$ (5)
The formulation of blur reconstruction loss is the same as in the original 3D-GS [13], it differs in utilizing blurry images as supervision and jointly optimizing the 3D-GS attributes and the camera trajectories, thus facilitating an initial deblurring reconstruction of 3D-GS.
Event Reconstruction Loss. Leveraging the abundant high-frequency information offered by the event streams, we further adopt an event reconstruction loss to aid in 3D-GS optimization. Specifically, we uniformly divide the exposure time into $m=n-1$ intervals, each with a duration of $\frac{\tau}{m}$. Subsequently, we integrate the recorded event stream along these time intervals using Eq. 1, resulting in $m$ event maps $\{\mathbf{E}_{i}\}_{i=1}^{m}$ to serve as event level supervision. During training, for the $j$-th blurry view, we convert the rendered image sequence $\{\mathbf{\widetilde{I}}_{i}\}_{i=1}^{n}$ on the camera trajectory into event maps $\{\mathbf{\widetilde{E}}_{i}\}_{i=1}^{m}$, using a differentiable event simulator [31, 9], and constrain the discrepancies between the simulated event maps and the ground truth event maps, expressed as:
$\mathcal{L}_{event}=\frac{1}{m}\sum_{i=1}^{m}\|\mathbf{E}_{i}-\mathbf{\widetilde{E}}_{i}\|_{1}.$ (6)
The event reconstruction loss further aids in recovering a sharp 3D-GS with improved texture details.
3.4 Event-assisted Geometry Regularization
The blurry color images are captured only during the exposure time and are much sparser than the event stream. Relying on such low-quality image-level supervision may cause the 3D-GS to overfit on the training images, resulting in significant floaters and inferior geometry, which affects the quality of novel view synthesis. Leveraging Eq. 1 and Eq. 4, given the continuously recorded event streams $\mathbf{E}(t)$, we can derive continuous grayscale intensity images $\mathbf{G}(t)$ that are rich in geometric information and can function beyond the exposure time. Motivated by this, we further propose two event-assisted geometry regularization terms to aid in 3D-GS training.
Intensity Reconstruction Loss. As shown in Figure. 2(C.1), during training, we randomly sample continuous time $t$ between the interval of two adjacent blurry image, and derive the grayscale intensity image $\mathbf{G}(t)$ using Eq. 2. We then minimize the difference between it and the rendered intensity image from 3D-GS, expressed as:
$\mathcal{L}_{int}=(1-\lambda_{2})\cdot\|\mathbf{G}(t)-\mathbf{\widetilde{G}}(t)\|_{1}+\lambda_{2}\cdot\text{D-SSIM}(\mathbf{G}(t),\mathbf{\widetilde{G}}(t)),$ (7)
where $\mathbf{\widetilde{G}}(t)$ is converted from the colored render result.
Intensity-aware Depth Regularization Loss. As shown in Figure. 2(C.2), to further improve the geometry of 3D-GS, inspired by [8, 4], we adopt an intensity-aware depth regularization loss during training, defined as:
$\mathcal{L}_{depth}=$ $\frac{1}{N}\sum_{x,y}(~{}|\partial_{x}\mathbf{\widetilde{D}}_{xy}(t)|e^{-\beta|\partial_{x}\mathbf{G}_{xy}(t)|}$ (8)
$+|\partial_{y}\mathbf{\widetilde{D}}_{xy}(t)|e^{-\beta|\partial_{y}\mathbf{G}_{xy}(t)|}~{}),$
where $\mathbf{\widetilde{D}}(t)$ is the rendered depth map, $(x,y)$ denotes the pixel location, $N$ is the total number of pixels, and $\beta$ is set to $2$ in our experiments. The horizontal and vertical gradients are calculated by applying convolution operations with $5\times 5$ Sobel kernels [39]. This regularization is founded on the observation that depth transitions in an image often correspond to changes in intensity. Therefore, it ensures that the spatial variation of depth closely matches that of the intensity image, thereby reducing geometric artifacts at object boundaries.
Table 1: Quantitative comparisons of novel view synthesis across large-scale, medium-scale, object-level, and real-world scenes. The table reports the average performance for each scale, demonstrating that our method consistently surpasses previous state-of-the-art approaches across all metrics. Best-performing results are highlighted in bold and second-best results in underline.
Scene Type Metric B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
The total loss function is the combination of the above losses, defined as:
$\mathcal{L}_{total}=$ $\lambda_{blur}\mathcal{L}_{blur}+\lambda_{event}\mathcal{L}_{event}$ (9)
$+\lambda_{int}\mathcal{L}_{int}+\lambda_{depth}\mathcal{L}_{depth}.$
4 Experiments
4.1 Implementation Details
Progressive Training. We implemented EvaGaussians based on the official code of 3D-GS [13]. The training process spans 50,000 iterations, with an event reconstruction loss introduced after a 3,000-iteration warmup and we omit the densification process to streamline and simplify the subsequent optimization. Additionally, we adopt a coarse-to-fine training strategy, starting with rendering at a low resolution (0.3$\times$ downsampling in the early 30% iterations) and progressively increasing the size of the rendered views to full resolution. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
Hyperparameter Setting. During the training process, we set $\lambda_{1}=0.2$, $\lambda_{blur}=1.0$, $\lambda_{depth}=1.0e^{-2}$, $\lambda_{event}=5.0e^{-3}$ and $\lambda_{int}=1.0e^{-3}$ for the loss function, and used $n=9$ for the number of poses to be optimized during the exposure time. In implementing the loss $\mathcal{L}_{event}$, we configured the positive threshold as ${c}_{pos}=0.25$ and the negative threshold as ${c}_{neg}=0.25$ for synthetic scenes, and set ${c}_{pos}=0.197$ and ${c}_{neg}=0.241$ for real scenes.
4.2 Datasets
To facilitate a comprehensive evaluation, we introduce two novel datasets, with an overview provided below. Detailed information is presented in the supplementary.
EvaGaussians-Blender Dataset. We construct a synthetic dataset covering a variety of scene scales, coupling with diverse camera trajectories and event data. For large-scale scenes, we employ Blender to craft five distinct scenes, including city blocks and natural landscapes. For medium-scale scenes, we craft three scenes using Blender, and redesign the camera trajectories of four scenes from DeblurNeRF [25]. For object-level scenes, we create six scenes based on the NeRF-synthetic [26] dataset. We simulate motion blur by manually placing multi-view cameras, randomly adjusting camera poses, and performing linear interpolation between the original and perturbed positions for each view. The images are rendered from these interpolated poses and blended in RGB space to produce the final blurry images. The corresponding event streams are simulated using ESIM [31] and V2E [9]. The resulting large-scale and medium-scale scenes comprise 35 views of blurry images along with their corresponding event data, whereas the object-level scenes feature 100 views of blurry images.
Figure 3: Qualitative comparison on the synthetic and real dataset. We show the rendering novel views on the top section (a) and exhibit both novel view synthesis results and input view deblurring results on the bottom section (b). It shows that our method achieves better performance in recovering the training blurry views as well as rendering novel views. More results are presented in the supplementary.
EvaGaussians-DAVIS Dataset. We manually recorded five real-world scenes using the Color DAVIS346 event camera [38], which has a resolution of $346\times 260$ pixels and an exposure time of 100 milliseconds for the RGB frames. The dataset includes three object-level scenes and two indoor scenes. After processing, the final dataset consists of 30 images per scene, along with the recorded event streams, each showcasing various blur and lighting conditions.
4.3 Experiment Settings
Baselines. We compare our method with three types of baselines: 1) NeRF [26] and 3D-GS [13] directly trained on the blurry images, referring to as B-NeRF and B-3DGS. 2) Deblur rendering methods, including BAD-NeRF [41], BAD-GS [52], ${\textnormal{E}^{2}}$eRF [30], and EDNeRF [3]. Among these, the first two methods simulate motion blur and optimize camera trajectories without event stream, whereas the latter two are event-assisted methods without camera trajectory optimization. 3) Image deblur methods, including UFP [6] (single-image deblurring), EDI [29] (event-based deblurring), and EFNet [35] (learnable event-based deblurring). We process input blurry images with them and train the vanilla 3D-GS with pre-deblurred images. The resulting baselines are referred to as UFP-GS, EDI-GS, and EFN-GS.
Evaluation Metrics. For synthetic datasets, we employ the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [42], and VGG-based Learned Perceptual Image Patch Similarity (LPIPS) [50] to evaluate the similarity between rendered novel views and ground-truth novel views. For real-world datasets, since the sharp ground-truth images are unavailable, we utilize several No-Reference Image Quality Assessment (NR-IQA) metrics for evaluation, including BRISQUE [27], NIQE [28], PIQE [40], RankIQA [21], and MetaIQA [53], which allow for image evaluation when lacking ground truth images.
4.4 Synthetic Data Experiments
We evaluate our approach across a variety of scenes, including large-scale scenes, medium-scale scenes, and object-level scenes. Quantitative assessments of novel view synthesis are shown in the first three rows of Table. 1. The deblurring results of input views are detailed in the supplementary. It can be found that our method achieves substantial improvements in most of the metrics, especially in challenging large scenes. Specifically, both B-NeRF and B-3DGS produce blurry novel views since they are directly trained on blurred images. The image deblurring-based baselines, UFP-GS, EDI-GS and EFN-GS, also produced inferior results, because the image deblurring process potentially corrupts the 3D consistency of the training images. Notably, our approach outperforms BAD-GS [52] and BAD-NeRF [41], due to their limited capability in modeling complex textures. In addition, our method also surpasses the event-assisted methods ${\textnormal{E}^{2}}$eRF [30] and EDNeRF [3] in producing high-quality novel views with intricate details, with better training and rendering efficiency. An extended analysis of all the baselines is provided in the supplementary.
The qualitative results are illustrated in Figure. 3, where the first three rows of Figure. 3(a) shows novel view synthesis results, and Figure. 3(b) shows both novel view and deblurring view synthesis results. More visualization results are provided in the supplementary. It can be found that although ${\textnormal{E}^{2}}$eRF [30] performs well in object-level scenes, it struggles in medium and large-scale scene modeling, producing significant blurring results. Additionally, BAD-GS [52] falls short in regions with significant color and depth variations, and produces overly smooth background textures. Although EDNeRF [3] exhibits overall satisfactory performance, its complex network architecture prolongs the training time (about 7 hours per scene) and precludes real-time rendering. In comparison, our method overcomes the baselines in producing high-fidelity novel views, and significantly reducing training time as well as demonstrating substantial advantages in real-time application scenarios.
Table 2: Quantitative ablation on proposed loss functions. Best-performing results are highlighted in bold and second results in underline.
Large-scale Medium-scale Real Scene
[…]
24.98 .7865 .2986 29.10 .8954 .1673 57.32 8.799 43.26 .1977 6.125
25.71 .7949 .2745 29.94 .9098 .1551 56.16 8.582 42.51 .2009 5.066
25.54 .8003 .2711 30.05 .9139 .1542 56.17 8.499 42.23 .1996 4.923
26.02 .8064 .2680 30.47 .9164 .1519 53.95 8.371 41.52 .2129 4.895
Table 3: Ablation study about the impact of pose optimization.
ATE $\downarrow$ Medium-scale Large-scale
Initial poses $0.0598\pm 0.0167$ $0.3825\pm 0.1245$
Optimized poses $0.0461\pm 0.0129$ $0.0489\pm 0.0112$
Table 4: Robustness against motion blur level.
Blur level Mild Blur Medium Blur Strong Blur
[…]
4.5 Real-world Data Experiments
We present the quantitative results on the captured real-world data in the last row of Table. 1. It can be found that our method achieves superior performance compared to other approaches. Specifically, for NR-IQA metrics, we achieve improvements in BRISQUE [27], NIQE [28], PIQE [40], and RankIQA [21] by 15.38%, 19.50%, 11.49%, and 22.83% respectively. We also achieve an increase in 19.38% in MetaIQA [53]. The qualitative comparisons are shown in the last row of Figure. 3(a) and in the supplementary, which further demonstrate that our method is capable of reconstructing detailed textures, ultimately achieving higher-quality novel view synthesis.
4.6 Ablation Study
Camera Poses Optimization. We firstly conduct ablations to investigate the effect of the number of camera poses optimized in the exposure time. We select five large scenes from our synthetic dataset for evaluation. In the experiments, we vary the number of camera poses, denoted as $n$, from 5, 9, 13, and 17. The quantitative results of the novel view rendering are displayed in Figure. 4. It indicates that the results reach a bottleneck at 9 poses. Beyond this point, the improvements are limited and may potentially lead to local convergence issues. Based on these experiments, we choose $n=9$ camera poses to achieve a balance between rendering performance and training efficiency. Here, we also provide comparison with BAD-NeRF [41] and BAD-GS [52]. These two methods typically use linear interpolation to obtain camera trajectory, while our camera trajectories are estimated from the decomposed latent images, which provides more accurate initialization and helps our method achieves better performance. Moreover, we conduct quantitative experiments using 9 camera poses to compute ATE (Average Trajectory Error) of the initial poses produced by COLMAP [34] and the optimized poses, the results are shown in Table. 3, which validates the effectiveness of pose optimization.
Effectiveness of The Loss Functions. We conduct novel view synthesis experiments on the proposed datasets to validate the effectiveness of the training losses. The quantitative results, as shown in Table. 2, indicate that using only the blur reconstruction loss leads to suboptimal outputs, performing poorly and lacking high-frequency details on both synthetic and real-world datasets. In contrast, incorporating $\mathcal{L}_{\text{event}}$, $\mathcal{L}_{\text{depth}}$, and $\mathcal{L}_{\text{int}}$ enables our proposed method to produce high-fidelity novel views with intricate details.
Robustness Against Motion Blur Levels. To validate the robustness of our method in handling different levels of motion blur, we set up three different camera speeds in the city blocks scene of the synthetic dataset to obtain images with varying degrees of blur. Images with mild blur are captured at half the default camera motion speed, images with medium blur are captured at the default motion speed, and images with strong blur are captured at twice the default motion speed. The quantitative results of novel view synthesis are listed in Table. 4. It can be observed that the results show no significant fluctuations across different levels of blur, demonstrating the robustness of our method to varying motion blur levels. Please refer to the supplementary for more ablations of our method.
Figure 4: Ablation on number of poses in the camera trajectory.
5 Conclusions
This paper introduces Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel framework that seamlessly integrates the event streams captured by an event camera into the training of 3D-GS, effectively addressing the challenges of reconstructing high-quality 3D-GS from motion-blurred images. We contribute two novel datasets and conduct comprehensive experiments. The results demonstrate that our method outperforms previous state-of-the-art deblurring rendering techniques in terms of novel view synthesis quality, without sacrificing inference efficiency. Despite its promising performance, our method may still face challenges when reconstructing scenes with extremely intricate textures from severely blurred images. We will release our code and dataset for future research.
Supplementary Material
Appendix A Dataset Details
A.1 EvaGaussians-Blender Dataset
A.1.1 Dataset Overview
We use Blender [5] to craft nine indoor and outdoor 3D scenes, and further incorporate four 3D scenes from DeblurNeRF [24] and six 3D objects from the NeRF-Synthetic dataset [26] as our base scenes. We then design various camera trajectories to simulate motion-blurred images on these base scenes, and generate the corresponding event streams using V2E [9]. Visualization of our crafted scenes are shown in Figure. 5 and Figure. 8, and an overview of the design purposes of these crafted scenes is provided below:
Figure 5: Visualization of EvaGaussians-Blender Indoor Scenes. The sizes of the Café and Classroom scenes are approximately ${15\times 7\times 4}$ meters, while the Dormitory scene is approximately ${5\times 5\times 4}$ meters (with an additional outdoor garden, making the overall scene size ${20\times 20\times 6}$ meters).
Figure 6: Visualization of Reprojection Errors and Epipolar Errors. The figure illustrates the 50 sets of reprojection errors and epipolar errors generated during the calibration process. The reprojection error $e_{r}$ represents the average discrepancy between the observed points and the projected points, calculated as shown in Eq. 10. The epipolar error $e_{\text{epipolar}}$ represents the average distance between points in one camera and the epipolar lines calculated from the other camera for each pair of images, calculated as shown in Eq. 11. As shown in the figure, the average reprojection error is approximately 0.5, and the average epipolar error is approximately 0.05, indicating a high level of accuracy in the calibration process.
Figure 7: Visualization of pose accuracy in different level of motion blur.
Figure 8: Visualization of EvaGaussians-Blender Outdoor Scenes. These scenes include rich details ane diverse components like sky, lake, river, desert, forest, cities, roads. All scenes cover an area of more than 1 square kilometer.
Indoor Scenes
  • Classroom: A typical classroom setting featuring desks, chairs, a blackboard, and educational posters. This scene is designed to simulate an academic environment, ideal for educational and surveillance applications.
  • Café: A cozy café with tables, chairs, a counter, and various decorations. This scene mimics a social setting, providing a dynamic backdrop for testing social interaction algorithms and retail analytics.
  • Dormitory: A student dormitory room equipped with beds, study desks, personal belongings, and typical dorm furniture. This scene represents a personal living space, useful for smart home and security applications.
Outdoor Scenes
  • Desert: A vast, arid landscape with sand dunes and sparse vegetation. This scene is perfect for testing navigation and object detection in harsh, unstructured environments.
  • City Blocks: Urban scenes featuring streets, buildings, vehicles, and pedestrians. This environment is essential for autonomous driving, urban planning, and smart city applications.
  • Lake: A serene natural setting with dense forests surrounding a tranquil lake. This scene provides a complex environment for testing outdoor navigation, environmental monitoring, and wildlife tracking.
  • Forests: A rugged terrain with forested areas and scattered boulders. This scene is useful for off-road navigation and geological survey applications.
  • Venice: A picturesque representation of Venice with canals, bridges, and historic architecture. This scene offers a unique setting for cultural heritage preservation, tourism, and urban analytics.
  • London: A bustling cityscape of London with iconic landmarks, streets, and a dynamic urban environment. This scene supports applications in tourism, traffic management, and city modeling.
Figure 9: Visualization of Camera Trajectory. The trajectories depicted were manually configured within Blender [5] to ensure precise control over the camera paths. For the purpose of visualization, these trajectories have been normalized.
A.1.2 Camera Settings
To render the base scenes and simulate motion blur, we configure the virtual camera in Blender with a resolution of $400\times 600$, and set the scaling factor to $1.0$. The virtual camera utilized a perspective model with a shutter speed of $1/180$ seconds. Subsequently, we developed a dedicated script to generate camera trajectory and motion blur. An example is shown in Figure. 9. Along each predefined virtual camera trajectory, we uniformly sampled 35 camera poses, adding a certain level of jitter to create the training set. We recorded the start and end time of the camera exposure time, the positions, and 20 intermediate frames during the exposure time (obtained through linear interpolation between the start and end positions). We then uniformly sample 100 camera poses along the same trajectory to form the test set. Using the event camera simulator from V2E [9], we simulate the event stream for each camera trajectory and synthesize the event bins from the event stream at the start and end of the exposure time.
A.2 EvaGaussians-DAVIS Dataset
We use the color DAVIS346 event camera [1] to record our real-world event and RGB sequences and utilize the default camera settings provided in the DV software that comes with the camera. We name the five captured scenes as desk & chair, washroom, pokémon, pillow, and bag.
A.2.1 Camera Calibration
Figure 10: Illustration of Camera Calibration. The left panel shows the checkerboard pattern captured from various positions and angles, with detected corner points utilized for calibration. The right panel presents the calibrated checkerboard pattern, demonstrating the corresponding points and lines between two cameras, which reflect the geometric relationship and accuracy achieved after calibration. Different colored lines indicate the correspondences between points during the calibration process.
We calibrated the event camera using the DV software provided by DAVIS. During the calibration process, we used a $6\times 9$ checkerboard pattern with a square size of 30 mm. In the software configuration, we set the width to 9, height to 6, and square size to 30 mm. We then ran the calibration module and moved the calibration pattern in front of the camera. The software detected the pattern and collected images, highlighting the detected area in green, as shown in Figure. 10. We set the minimum detections parameter to 50 to ensure a sufficient number of samples and used the consecutive detections parameter to ensure consistent pattern detection. Additionally, we enabled the image verification option to check the collected images in real-time, discarding inaccurately detected images and replacing them with new ones. We evaluate the calibration accuracy using the reprojection error $e_{r}$ as Eq. 10 and the epipolar error $e_{\text{epipolar}}$ as Eq. 11 in stereo calibration. The reprojection error is calculated as follows:
$e_{r}=\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-\hat{x}_{i}\|$ (10)
where $x_{i}$ represents the observed points and $\hat{x}_{i}$ represents the projected points. The epipolar error is calculated as the average epipolar error for each point in all collected images. For each pair of images, the error is calculated as the sum of the distances between the points in one camera and the epipolar lines calculated from the other camera ($m$ is the number of acquired images, $n$ is the number of points). The formula is as follows:
$e_{\text{epipolar}}=\frac{1}{m\times n}\sum_{i=1}^{m}\sum_{j=1}^{n}\left[d(P1_{i,j},l_{2,i,j})+d(P2_{i,j},l_{1,i,j})\right]$ (11)
where $P1_{i,j}$ and $P2_{i,j}$ are the projection points of the $j$th point in the $i$th image in two cameras, and $l_{1,i,j}$ and $l_{2,i,j}$ are the epipolar lines corresponding to the $j$th point in the $i$th image calculated from the other camera. The maximum allowable error can be set under Max Reprojection Error. The stereo calibration also calculates the error caused by the epipolar constraint, which can be set under Max Epipolar Error. Once the calibration is successful, the results are saved and the undistorted output is displayed. The visualization results of the two types of errors are shown in Figure. 6. This process ensures the accuracy of the calibration, thereby improving the measurement accuracy and stability in subsequent applications.
A.2.2 Camera Settings
We recorded the five real scenes using the calibration parameters obtained during the calibration process. By adjusting the indoor lighting and shooting angles, we ensured the richness of the recorded scene details. The adopted event camera has a spatial resolution of $346\times 260$, a temporal resolution of 1 $\mu$s, a typical latency of less than 1 ms, a maximum throughput of 12 MEps, and a dynamic range of approximately 120 $dB$ (with 50% of the pixels responding to 80% contrast changes under 0.1-100k lux conditions). The contrast sensitivity is 14.3% (ON) and 22.5% (OFF) (with 50% of the pixels responding). These parameters ensure that the event camera can stably and efficiently record scene information under various lighting conditions and dynamic ranges.
Table 5: Quantitative comparisons of DVS on object-level scenes. The results indicate that our method outperforms previous state-of-the-art approaches, consistently achieving better performance across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
Table 6: Quantitative comparisons of DVS on the medium-scale scenes. The results show that our method surpasses previous state-of-the-art approaches, achieving better performance consistently across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
Table 7: Quantitative comparison of DVS on large-scale scenes. The results demonstrate that our method consistently achieves better performance across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
Table 8: The Novel View Synthesis Results of PSNR $\uparrow$ in the EvaGaussians-Blender Dataset. The highest values in each category are highlighted in bold to indicate the best results.
Models Medium Large
Classroom Dormitory Café Pool Cozyroom Factory Tanabata Desert City Blocks London Forests Lake
Blurry-NeRF 25.42 26.72 21.23 26.66 15.42 26.76 26.38 19.04 19.34 20.00 23.60 24.67
Blurry-GS 25.59 26.83 21.34 26.17 20.76 26.48 26.46 19.67 19.48 20.18 23.75 24.34
UFP-GS 28.64 29.68 21.54 26.79 21.62 27.60 28.77 20.78 20.85 16.61 22.96 25.60
EDI-GS 28.49 29.90 22.62 26.36 21.17 29.06 27.46 20.66 20.68 20.27 23.97 25.99
EFNET-GS 28.59 29.48 22.04 26.36 21.42 27.57 27.43 20.62 21.70 20.36 25.57 25.18
$\textnormal{E}^{2}$eRF 28.87 30.77 26.16 28.92 21.23 29.85 28.65 20.02 21.78 21.30 25.90 25.78
BAD-NeRF 30.28 31.23 27.18 28.71 21.68 30.28 29.85 21.09 21.91 22.91 26.35 26.99
BAD-Gaussians 30.28 31.23 27.18 28.72 21.68 30.28 29.85 21.10 21.93 22.93 26.35 26.98
EvDeblurNeRF 31.83 28.95 28.66 29.69 22.01 31.20 30.02 21.62 22.23 23.88 27.10 28.29
Ours 34.38 32.97 30.41 30.26 22.71 31.85 30.71 24.88 23.71 23.99 27.62 29.90
Table 9: The Novel View Synthesis of SSIM $\uparrow$ in the EvaGaussians-Blender Dataset. The highest values in each category are highlighted in bold to indicate the best results.
Models Medium Large
Classroom Dormitory Café Pool Cozyroom Factory Tanabata Desert City Blocks London Forests Lake
Blurry-NeRF 0.7086 0.8281 0.5682 0.7442 0.5098 0.8567 0.8057 0.6023 0.6325 0.6732 0.7002 0.7823
Blurry-GS 0.7154 0.8312 0.5638 0.7265 0.7632 0.8519 0.8064 0.6386 0.6341 0.6819 0.7026 0.7807
UFP-GS 0.8701 0.9281 0.5706 0.7527 0.7729 0.8640 0.8569 0.6172 0.6144 0.5807 0.6911 0.7968
EDI-GS 0.8456 0.9291 0.5977 0.7288 0.7703 0.8963 0.8408 0.6567 0.5655 0.6882 0.7158 0.8012
EFNET-GS 0.8692 0.9259 0.5841 0.7288 0.7797 0.8643 0.8345 0.6164 0.6366 0.6069 0.7628 0.7901
$\textnormal{E}^{2}$eRF 0.8723 0.9319 0.7869 0.8795 0.7724 0.9245 0.8915 0.6169 0.6822 0.6734 0.7467 0.8139
BAD-NeRF 0.8978 0.9337 0.8041 0.8794 0.7764 0.9353 0.9271 0.6314 0.6867 0.6932 0.7583 0.8919
BAD-Gaussians 0.8992 0.9351 0.8031 0.8781 0.7755 0.9345 0.9266 0.6346 0.6851 0.6944 0.7578 0.8908
EvDeblurNeRF 0.9023 0.8935 0.8513 0.8885 0.7854 0.9454 0.9311 0.6589 0.7023 0.7159 0.7726 0.9129
Ours 0.9402 0.9580 0.9033 0.9108 0.8052 0.9592 0.9382 0.8152 0.7405 0.7145 0.8152 0.9465
Table 10: The Novel View Synthesis of LPIPS $\downarrow$ in the EvaGaussians-Blender Dataset. The highest values in each category are highlighted in bold to indicate the best results.
Models Medium Large
Classroom Dormitory Café Pool Cozyroom Factory Tanabata Desert City Blocks London Forests Lake
Blurry-NeRF 0.3987 0.2998 0.4528 0.3848 0.5545 0.2063 0.2348 0.4447 0.5273 0.3971 0.3425 0.4127
Blurry-GS 0.3824 0.2873 0.4554 0.4198 0.2432 0.2091 0.2335 0.4231 0.4116 0.3925 0.3379 0.4206
UFP-GS 0.2838 0.1361 0.4511 0.3646 0.2331 0.1928 0.1856 0.3816 0.3342 0.4069 0.3383 0.4069
EDI-GS 0.2914 0.1267 0.4038 0.4102 0.2388 0.1446 0.1911 0.4062 0.3857 0.3901 0.3295 0.4001
EFNET-GS 0.2846 0.1373 0.4529 0.4102 0.2345 0.1926 0.1958 0.3912 0.3222 0.3932 0.2962 0.4128
$\textnormal{E}^{2}$eRF 0.2821 0.1165 0.2871 0.2048 0.2369 0.1073 0.1546 0.3938 0.3713 0.3928 0.3589 0.3588
BAD-NeRF 0.2365 0.1083 0.2715 0.2103 0.2278 0.0992 0.1224 0.3947 0.3618 0.3224 0.3467 0.3142
BAD-Gaussians 0.2384 0.1078 0.2695 0.2094 0.2262 0.0985 0.1215 0.3965 0.3604 0.3215 0.3452 0.3129
EvDeblurNeRF 0.2217 0.1455 0.2447 0.1926 0.1942 0.0768 0.1086 0.3823 0.3497 0.2972 0.3163 0.2941
Ours 0.1927 0.0928 0.2311 0.1859 0.1873 0.0715 0.1023 0.2053 0.2835 0.2983 0.2857 0.2674
Table 11: The Novel View Synthesis of BRISQUE in the EvaGaussians-DAVIS Dataset. The highest values in each category are highlighted in bold to indicate the best results.
      Models       BRISQUE $\downarrow$
      Desk & Chair       Washroom       Pokémon       Pillow       Bag       Average
      B-NeRF       63.9428       102.7828       109.2711       97.4778       87.7699       92.2489
      B-3DGS       51.1542       82.2262       87.4169       77.9823       70.2159       73.7991
      UFP-GS       43.6932       69.7354       74.6684       66.6098       59.9711       62.9356
      EDI-GS       43.4811       69.9923       74.3044       66.2849       59.6835       62.7492
      EFN-GS       43.5235       69.8473       74.2875       66.3158       59.6523       62.7253
      ${\textnormal{E}^{2}}$eRF       36.2148       64.4332       77.0112       67.1324       62.8063       61.5196
      BAD-NeRF       42.6285       68.5219       72.8474       64.9852       58.5133       61.4993
      BAD-GS       42.2065       67.8434       72.1261       64.3418       57.9339       60.8903
      EDNeRF       34.5687       61.5044       73.5012       64.0809       59.4969       58.6304
      Ours       32.9225       58.5756       70.0211       61.0294       58.1876       56.1472
Table 12: Robustness of pose optimization.
           ATE $\downarrow$            Mild Blur            Medium Blur            Strong Blur
           COLMAP poses            $0.0598\pm 0.0167$            $0.0783\pm 0.0245$            $0.0973\pm 0.0352$
           Optimized poses            $0.0461\pm 0.0129$            $0.0586\pm 0.0149$            $0.0677\pm 0.0208$
Table 13: The Novel View Synthesis of NIQE in the EvaGaussians-DAVIS Dataset.
      Models       NIQE $\downarrow$
      Desk & Chair       Washroom       Pokémon       Pillow       Bag       Average
      B-NeRF       11.4380       15.5386       16.9153       16.2316       14.8844       15.0016
      B-3DGS       9.1504       12.4309       13.5323       12.9853       11.9075       12.0113
      UFP-GS       7.7736       10.6170       11.3827       10.9266       10.1715       10.1743
      EDI-GS       7.7778       10.5662       11.5024       11.0375       10.1214       10.2011
      EFN-GS       7.8235       10.5412       11.4868       11.0743       10.1288       10.2109
      ${\textnormal{E}^{2}}$eRF       6.8907       9.0924       11.6383       10.1830       9.3954       9.4400
      BAD-NeRF       7.6253       10.3590       11.2769       10.8211       9.9229       10.0011
      BAD-GS       7.5498       10.2565       11.1652       10.7139       9.8247       9.9020
      EDNeRF       6.5775       8.6791       11.1092       9.7201       8.9684       9.0109
      Ours       6.1898       8.1117       10.2662       8.9455       8.3421       8.3711
Table 14: The Novel View Synthesis of PIQE in the EvaGaussians-DAVIS Dataset.
      Models       PIQE $\downarrow$
      Desk & Chair       Washroom       Pokémon       Pillow       Bag       Average
      B-NeRF       58.4566       68.8701       73.4701       54.0862       74.7257       65.9217
      B-3DGS       46.7653       55.0961       58.7761       43.2689       59.7806       52.7374
      UFP-GS       39.9453       47.0612       50.2026       36.9659       50.9631       45.0276
      EDI-GS       39.7505       46.8317       49.9597       36.7786       50.8135       44.8268
      EFN-GS       39.7748       46.7951       49.9346       36.7891       50.8275       44.8242
      ${\textnormal{E}^{2}}$eRF       40.9825       49.2988       50.4868       39.7896       53.2203       46.7556
      BAD-NeRF       38.9711       45.9134       48.9801       36.0574       49.8171       43.9478
      BAD-GS       38.5852       45.4588       48.4951       35.7004       49.3239       43.5127
      EDNeRF       39.1197       47.0580       48.1919       37.9810       50.8012       44.6304
      Ours       36.5717       44.3304       44.0641       35.3059       47.3749       41.5294
Table 15: The Novel View Synthesis of MetaIQA in the EvaGaussians-DAVIS Dataset.
      Models       MetaIQA $\uparrow$
      Desk & Chair       Washroom       Pokémon       Pillow       Bag       Average
      B-NeRF       0.1419       0.1272       0.1085       0.1122       0.1307       0.1241
      B-3DGS       0.1621       0.1454       0.1240       0.1283       0.1494       0.1418
      UFP-GS       0.1976       0.1774       0.1521       0.1564       0.1823       0.1732
      EDI-GS       0.1986       0.1780       0.1518       0.1571       0.1830       0.1737
      EFN-GS       0.1928       0.1835       0.1579       0.1643       0.1799       0.1757
      ${\textnormal{E}^{2}}$eRF       0.1959       0.2020       0.1619       0.1723       0.1722       0.1809
      BAD-NeRF       0.2027       0.1817       0.1549       0.1603       0.1867       0.1773
      BAD-GS       0.2047       0.1835       0.1565       0.1620       0.1886       0.1790
      EDNeRF       0.2067       0.2132       0.1709       0.1819       0.1817       0.1909
      Ours       0.2135       0.2188       0.1975       0.2188       0.2160       0.2129
Table 16: The Novel View Synthesis of RankIQA in the EvaGaussians-DAVIS Dataset.
      Models       RankIQA $\downarrow$
      Desk & Chair       Washroom       Pokémon       Pillow       Bag       Average
      B-NeRF       7.4896       10.4454       10.7921       9.6578       8.7541       9.4278
      B-3DGS       5.9917       8.3563       8.6337       7.7262       7.0032       7.5422
      UFP-GS       5.1153       7.1215       7.3741       6.6005       5.9829       6.4389
      EDI-GS       5.0929       7.1029       7.3386       6.5673       5.9528       6.4109
      EFN-GS       5.1046       7.0893       7.3158       6.5789       5.9682       6.4114
      ${\textnormal{E}^{2}}$eRF       4.5141       6.4997       5.7764       5.7600       5.3166       5.5733
      BAD-NeRF       4.9931       6.9636       7.1948       6.4385       5.8360       6.2852
      BAD-GS       4.9436       6.8947       7.1235       6.3748       5.7783       6.2230
      EDNeRF       4.3089       6.2043       5.5138       5.4981       5.0749       5.3200
      Ours       3.9369       5.6564       5.0867       5.1538       4.6422       4.8952
Figure 11: NVS results on the EvaGaussian-DAVIS dataset. The first column shows the blurry image used for training, and the following rows show the deblurring results of different methods. The results demonstrate that our method consistently excels in reconstructing fine details compared to other methods [30, 52, 3].
Figure 12: Visualization of DVS and NVS of Object-level Scenes in the EvaGaussian-Blender Dataset. The DVS results are highlighted in the red bouding box. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods [26, 13, 29, 35, 30, 52, 3, 6].
Figure 13: Visualization of DVS and NVS of results. The DVS results are highlighted in the red bouding box. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods.
Figure 14: Visualization of Novel View Synthesis of All-redesigned Scenes with B-NeRF, B-3DGS, EDI-GS, EFN-GS and UFP-GS in the EvaGaussian-Blender Dataset.
Figure 15: Visualization of Novel View Synthesis of All-redesigned Scenes with ${\textnormal{E}^{2}}$eRF, BAD-GS, EDNeRF and EvAGS in the EvaGaussian-Blender Dataset.
Appendix B Detailed Experiments
B.1 Synthetic Data Experiments
B.1.1 Deblurring View Synthesis Comparison
Refering to [41], we additionally provide deblurring view synthesis (DVS) results on our proposed EvaGaussians-Blender dataset, and show more qualitative results of novel view synthesis (NVS). For object-level scenes, Table. 5 and Figure. 12 present the quantitative and qualitative results of ours and the comparison baselines across six synthetic scene sequences. From the qualitative results, it is evident that our method excels in reconstructing fine details and maintaining high fidelity in both NVS and DVS. In terms of quantitative results, our method outperforms baseline methods in most scenes. For medium-scale scenes and large-scale scenes, the quantitative results are shown in Table. 6 and Table. 7, and the qualitative results of NVS and DVS are shown in Figure. 13, which demonstrate that our model achieves better performance in both tasks.
B.1.2 Per-scene Comparison for Novel View Synthesis
In this subsection, we present a detailed per-scene analysis of the novel view synthesis performance in medium and large scale scenes from the EvaGaussians-Blender dataset, to evaluate the effectiveness of our method across different challenging scenes.
Table. 8 shows the PSNR value of the NVS results, which demonstrates that our proposed method consistently outperforms other approaches across various scenes. The detailed metrics for SSIM and LPIPS in Table. 9 and Table. 10 further show that our model excels in maintaining structural integrity and perceptual quality in synthesized views. Specifically, in medium-scale scenes, our method exhibits robust performance, particularly in complex environments where maintaining detail and minimizing artifacts are challenging. This can be demonstrated in scenes such as cozyroom and factory, where our method achieves significant improvements in both PSNR and SSIM. For large-scale scenes, scenes like desert and city blocks highlight the model’s capability to generalize across different scales and provide high-quality novel view synthesis.
We also present more qualitative results of novel view synthesis in Figure. 14 and Figure. 15. The results highlight our model’s ability to reconstruct fine details and maintain high color accuracy beyond the comparison baselines.
B.2 Real-world Data Experiments
In this section, we present a comprehensive per-scene analysis of NVS results on the EvaGaussians-DAVIS dataset. The qualitative results are shown in Figure. 11, where the first column shows the blurry image used for training, and the following rows show the deblur results of different methods. The results demonstrate that our method consistently excels in reconstructing fine details compared to other methods.
We further report the per-scene quantitative results to validate our robustness to different scenes. As introudced in the main text, the adopted metrics include BRISQUE, NIQE, PIQE, MetaIQA, and RankIQA, which can effectively assess the quality of synthesized views in a no-reference manner. As shown in Table. 11, our model achieves the best BRISQUE scores across all scenes, highlighting its ability to produce visually appealing and less distorted images. For NIQE, as presented in Table. 13, our approach significantly outperforms the baselines, achieving the lowest average NIQE score. This demonstrates our method’s robustness in generating high-quality images with minimal perceptual artifacts. In terms of PIQE, Table. 14 shows that our model again leads in performance, achieving the lowest PIQE scores, which underscores the effectiveness of our model in preserving image details and reducing noise. Furthermore, our method excels in MetaIQA and RankIQA evaluations, as detailed in Tables. 15 and Table. 16, respectively. The highest MetaIQA scores and lowest RankIQA scores across most scenes affirm the overall better visual quality and fidelity of our synthesized views compared to baseline models. Overall, these results demonstrate the robustness of our method, particularly in handling complex scenes and maintaining high visual quality across diverse scenarios.
B.3 Ablation Study
In this section, we present an additional ablation study on the robustness of pose optimization in different blur level. We redesign three different levels of motion blur sequences in medium-scale scenes and compare the Average Trajectory Error (ATE) between the initial poses produced by COLMAP and the optimized poses. Figure. 7 illustrates the visualization of COLMAP poses and the optimized poses on the City Blocks scene. As the motion blur becomes more severe, the accuracy of the COLMAP poses is significantly impacted, while the optimized poses maintain a higher level of accuracy. In a horizontal comparison, the optimized poses better match the ground truth across various levels of blur, demonstrating the effectiveness of pose optimization. Table. 12 presents the quantitative results, further demonstrating the effectiveness and robustness of pose optimization in handling different levels of motion blur.
Appendix C Broader Impacts
Our proposed EvaGaussians leverages event cameras to assist novel view synthesis from low-quality, blurred images. It has the potential to bring about both positive and negative societal impacts.
On the positive side, our method can improve the efficiency of surveillance systems by reconstructing clear 3D images from low-quality footage, enabling better identification of individuals and objects in challenging conditions. This can bolster public safety and aid in criminal investigations. Additionally, the ability to reconstruct scenes from blurred inputs can enhance the performance of autonomous vehicles, drones, and robots, enabling them to navigate more accurately in poor visibility conditions, leading to safer and more efficient transportation and logistics. In situations where traditional cameras may struggle to capture clear images under extreme conditions, our method can provide valuable information for first responders and rescue teams, helping them make informed decisions and potentially saving lives. Furthermore, our technique can be applied to medical imaging, allowing for better visualization of internal structures and more accurate diagnoses, ultimately leading to improved patient outcomes.
On the negative side, the enhanced surveillance capabilities enabled by our method may raise privacy concerns. For example, our method could be used for malicious purposes, such as stalking or spying on individuals without their consent. It is important to establish regulations and guidelines to prevent such misuse.